OpenStudy (anonymous):

The revenue from selling x units of a product is given by y=-0.0002x^2+20x. how many units must be sold in order to have the greatest evenue?

OpenStudy (anonymous):

greatest will be at the vertex

OpenStudy (anonymous):

first coordinate of the vertex is \(-\frac{b}{2a}\) which in your case is \[-\frac{20}{2\times (-0.0002)}\]